OpenStudy (anonymous):

(-5p to the fourth power z to the sixth power u) 3 hint: apply the 3 exponent to each part..... I am lost please help!!!

OpenStudy (anonymous):

Assuming you are starting from the following: \[\left(\begin{matrix}-5p ^{4}z ^{6} \\ \end{matrix}\right)^{3}\] You would apply the power rule of exponents which states that if you have an exponent raised to an exponent you muplifply the two exponents to get a new exponent. For example: \[\left( a ^{2}b ^{3} \right)^{3}=a ^{2*3}b ^{3*3}=a ^{6}b ^{9}\]

OpenStudy (anonymous):

yes there is a u after the 6 though and I just don't understand it